# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v1

@@ 5 Conclusion
In this work, we presented a residual-based Chebyshev filtered subspace iteration (R-ChFSI) method tailored for large sparse Hermitian eigenvalue problems while being tolerant to inexact matrix-vector products. We have demonstrated that the proposed method significantly outperforms the standard Chebyshev filtered subspace iteration (ChFSI) approach through extensive analysis and numerical experiments. We have studied the performance and accuracy of R-ChFSI under different computational precisions, demonstrating that it maintains robust convergence properties even when using lower-precision arithmetic. This characteristic makes it particularly suitable for modern hardware architectures that prioritize low-precision operations for efficiency gains in AI training. The results indicate that R-ChFSI achieves similar residual tolerances as full-precision ChFSI while retaining the performance benefits of reduced precision arithmetic. Furthermore,
@@ 6 Acknowledgements
P.M. acknowledges Prathu Tiwari at Nvidia, Bangalore, for helping us run the benchmarks on NVIDIA DGX SuperPOD Eos system. The research also used the resources of PARAM Pravega at the Indian Institute of Science, supported by the National Supercomputing Mission (NSM) for testing few of the benchmarks on CPUs. N.K and K.R would like to acknowledge the Prime Minister Research Fellowship(PMRF) from the Ministry of Education India for financial support. P.M. acknowledges the Google India Research Award 2023 for financial support during the course of this work.

### v2

@@ 5 Conclusion
In this work, we presented a residual-based Chebyshev filtered subspace iteration (R-ChFSI) method tailored for large sparse Hermitian eigenvalue problems while being tolerant to inexact matrix-vector products. We have demonstrated that the proposed method significantly outperforms the standard Chebyshev filtered subspace iteration (ChFSI) approach through extensive analysis and numerical experiments. We have provided a mathematical justification demonstrating that this reformulation reduces the numerical error in the Chebyshev filtered subspace construction while employing inexact matrix-vector products compared to the traditional ChFSI recurrence relation. We have studied the performance and accuracy of R-ChFSI under different computational precisions, demonstrating that it maintains robust convergence properties even when using lower-precision arithmetic. This characteristic makes it particularly suitable for modern hardware architectures that prioritize low-precision operations for efficiency gains in AI training. The results indicate that R-ChFSI achieves similar residual tolerances as full-precision ChFSI while retaining the performance benefits of reduced precision arithmetic.
For standard eigenvalue problems, we demonstrate that the proposed R-ChFSI method leverages the advantages of low-precision arithmetic. This capability is critical in today’s computational landscape, where the use of low-precision arithmetic (e.g., FP32, TF32) on modern hardware accelerators can dramatically reduce computational costs while the proposed R-ChFSI method still allows us to achieve high accuracy. Our benchmarks demonstrate that, for large-scale eigenproblems, R-ChFSI attains residual norms in the range of $10^{-12}$ to $10^{-14}$ which is a significant improvement over traditional ChFSI methods, that stagnate at residuals around $10^{-6}$ when using lower precision. In the context of generalized eigenvalue problems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$, the proposed R-ChFSI method is particularly impactful due to its ability to incorporate approximate inverses of B. By replacing the exact computation of matrix inverses with computationally inexpensive approximate inverses (e.g., diagonal or block-diagonal approximations of B), the method not only reduces the overall computational burden but also maintains convergence robustness. This results in substantial performance gains, as the proposed approach achieves residual tolerances that are up to ten orders of magnitude lower than those obtainable with traditional ChFSI when solving generalized problems. We further note that even in the context of generalized eigenvalue problems, the proposed R-ChFSI method allows us to utilize low-precision arithmetic alongside approximate inverses, further enhancing the computational gains.
@@ 6 Acknowledgements
The authors gratefully acknowledge the seed grant from the Indian Institute of Science and the SERB Startup Research Grant from the Department of Science and Technology India (Grant Number: SRG/2020/002194). P.M. acknowledges Prathu Tiwari at Nvidia, Bangalore, for helping us run the benchmarks on NVIDIA DGX SuperPOD Eos system. The research also used the resources of PARAM Pravega at the Indian Institute of Science, supported by the National Supercomputing Mission (NSM) for testing few of the benchmarks on CPUs. N.K and K.R would like to acknowledge the Prime Minister Research Fellowship(PMRF) from the Ministry of Education India for financial support. P.M. acknowledges the Google India Research Award 2023 for financial support during the course of this work.